OpenStudy (anonymous):

If A can do a job in 8 days and b can do the same job in 12 days, how long would it take them both to complete the job working together? 1. 3 2. 4.8 3. 10 4. 5.8 5. 6

OpenStudy (anonymous):

You divide the product of the numbers by their sum. \[\frac{(8)(12)}{8+12}\]

OpenStudy (anonymous):

number 2.

jimthompson5910 (jim_thompson5910):

A can do the job in 8 days. So his rate is 1/8 of a job per day. B can do the job in 12 days, so his rate is 1/12 of a job per day. Add the two rates to get the total rate of 1/8+1/12 = 5/24. Now use the idea that Rate * Time = 1 where 1 represents 1 job completed. So we get (5/24)*t = 1 t = 24/5 t = 4 4/5 t = 4.8 So it will take 4.8 days for them to work together.